OpenStudy (anonymous):

Do anyone know the annual rate and time to double equation for an investment

OpenStudy (anonymous):

72/rate

OpenStudy (anonymous):

you need one or the other. rate or time. what exactly is the question? oh right, rule of 72 is a good approximation.

OpenStudy (anonymous):

\[Number of years 2 double = \frac{72}{Interest Rate} years\]

OpenStudy (anonymous):

Initial Investment $600 Amount after 10 years 19,205.00 What is the annual % rate? What is the time to double?

OpenStudy (anonymous):

here rate is equal to 41.425 %

OpenStudy (anonymous):

and time to double is 1.73 years

OpenStudy (anonymous):

just solve 2=(1+rate)^x

OpenStudy (anonymous):

solve \[600(1+r)^{10}=19,205\] for r \[(1+r)^{10}=\frac{3841}{120}=32\](rounded) \[1+r=\sqrt[10]{32}=\sqrt{2}\]

OpenStudy (anonymous):

that is a hell of a rate of interest

OpenStudy (anonymous):

\[r=\sqrt{2}-1=.4142\] as sheg said. but rounded it is \[.412421\] not \[.412425\]

OpenStudy (anonymous):

time to double solve \[1.41421^x=2\] \[x=\frac{\ln(2)}{\ln(1.4121)}\]

OpenStudy (phi):

I think sheg's interest rate assumes continuous compound interest and satellites is compounded yearly. don't understand sheg's doubling time. It should be very close to 2 years.

OpenStudy (anonymous):

@PHI dear friend in case of compound interest u earn interest on principal as well as interest earned formula for compounding is \[Amount = Principal \times (1 + r )^n\] now do this problem Principal,P = $ 1000 rate, r = 41.425 % Number of years, n = 2 years and see how much amount u have after 2 years

OpenStudy (anonymous):

http://www.moneychimp.com/features/rule72.htm

OpenStudy (anonymous):

actually if u will make aproximation then it will double in 1.73 yrs because i was using the rule of 72